Oneofus:

Find the difference in the volume and total area of a cylinder with both a radius and height of 1. r = 1, h = 1 The number of sq units of the total area exceeds the number of cu. units in the volume by 3 2

Oneofus:

3 pi?

Hero:

@Oneofus The Total Surface Area of a cylinder is represented by \(TSA = 2\pi r(r + h)\) The Volume of a Cylinder is represented by \(V = \pi r^2 h\) If \(r = 1, h = 1\) then: \(TSA = 2 \pi (1)(1 + 1) = 4 \pi\) \(V = \pi (1)^2 (1) = 3\pi\) So the difference between the Total Surface Area of a Cylinder and the Volume of a Cylinder is represented by: \(TSA - V\) in square units

Oneofus:

Thank you!

Hero:

Whoops, sorry \(V = \pi\) not \(3 \pi\)

Hero:

I had "3" in my head when I wrote that.

Oneofus:

o.o

Oneofus:

So its just a pi?

Hero:

But yes, 3 is correct as far as the answer to the question as you have correctly stated.

Oneofus:

alright, thank you

Hero:

I just made the clarification as to the value of the Volume

Hero:

\(V = \pi\) square units